there are three consecutive integers half the sum of the first and third is 52 what are the integers
Answer by checkley77(12569)

Let x, x+1 & x+2 be the 3 integers.
(x+x+2)/2+3=52
(2x+2)/2=52-3
(2x+2)/2=49
2x+2=2*49
2x+2=98
2x=98-2
2x=96
x=96/2
x=48 ans. for the smallest integer.
Proof:
(48+48+2)/2+3=52
(96+2)/2+3=52
98/2+3=52
49+3=52
52=52